OpenStudy (anonymous):

Write an equation in Slope-intercept form for the line that is parallel to the line y = -5/2 x + 6 and that passes through the given point (-4,-6). Show your work. Please help!!!

jimthompson5910 (jim_thompson5910):

you are given this equation \[\Large y = -\frac{5}{2}x + 6\] what is the slope of this given line?

OpenStudy (anonymous):

-5/2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

we need the equation of some line parallel to this one

jimthompson5910 (jim_thompson5910):

all parallel lines have the same slope so we know \[\Large m = -\frac{5}{2}\] for this new line

jimthompson5910 (jim_thompson5910):

we are told that this parallel line "passes through the given point (-4,-6)" all points are of the form (x,y) so (x,y) = (-4,-6) which means x = -4 y = -6

jimthompson5910 (jim_thompson5910):

now that we have the values of m, x and y, we can plug them into y = mx+b and solve for b

jimthompson5910 (jim_thompson5910):

\[\Large y = mx+b\] \[\Large y = -\frac{5}{2}x+b\] \[\Large -6 = -\frac{5}{2}(-4)+b\] solve for b

OpenStudy (anonymous):

which would be

jimthompson5910 (jim_thompson5910):

you tell me. It might be helpful to compute -5/2 times -4 to get \[\Large -\frac{5}{2}(-4) = \frac{-5}{2}*\frac{-4}{1}\] \[\Large -\frac{5}{2}(-4) = \frac{-5*(-4)}{2*1}\] \[\Large -\frac{5}{2}(-4) = \frac{20}{2}\] \[\Large -\frac{5}{2}(-4) = 10\]

jimthompson5910 (jim_thompson5910):

so, \[\Large -6 = -\frac{5}{2}(-4)+b\] turns into \[\Large -6 = 10+b\]

jimthompson5910 (jim_thompson5910):

so b = ??